A rectangular fish tank measures 30 cm by 70 cm by 40 cm high. The tank is filled with water to a depth of 37 cm.
What is the pressure at the bottom due to the water?
What is the total force of the water on the bottom?
1
Expert's answer
2012-03-30T11:29:02-0400
The preassure of the water atthe bottom is equal to
p = rho g h,
where

rho = 1000 kg/m^3 is the density of water,
g = 9.8 m/c^2
h =
0.37 cm is the height of water.

Hence

p = 1000 * 9.8 * 0.37 =
3626 Pa


The forse is equal to
F = rho g V
where
V = 0.3
* 0.7 * 0.37 = 0.0777 m^3.

So,

F = 1000 * 9.8 * 0.0777 = 761.46 N